# Balanced Diffusion-Guided Fusion for Multimodal Remote Sensing Classification

### v1

@@ Abstract
Deep learning-based techniques for the analysis of multimodal remote sensing data have become popular due to their ability to effectively integrate complementary spatial, spectral, and structural information from different sensors. Recently, denoising diffusion probabilistic models (DDPMs) have attracted attention in the remote sensing community due to their powerful ability to capture robust and complex spatial-spectral distributions. However, pre-training multimodal DDPMs may result in modality imbalance, and effectively leveraging diffusion features to guide complementary diversity feature extraction remains an open question. To address these issues, this paper proposes a balanced diffusion-guided fusion (BDGF) framework that leverages multimodal diffusion features to guide a multi-branch network for land-cover classification. Specifically, we propose an adaptive modality masking strategy to encourage the DDPMs to obtain a modality-balanced rather than spectral image-dominated data distribution. Subsequently, these diffusion features hierarchically guide feature extraction among CNN, Mamba, and transformer networks by integrating feature fusion, group channel attention, and cross-attention mechanisms. Finally, a mutual learning strategy is developed to enhance inter-branch collaboration by aligning the probability entropy and feature similarity of individual subnetworks. Extensive experiments on four multimodal remote sensing datasets demonstrate that the proposed method achieves superior classification performance. The code is available at https://github.com/HaoLiu-XDU/BDGF.
@@
Multimodal remote sensing data typically comprises images from various sensors, including hyperspectral images (HSI), synthetic aperture radar (SAR), and light detection and ranging (LiDAR). By fusing complementary spatial, spectral, and structural information, multimodal data facilitates an accurate classification [1, 2], thereby supporting applications in environmental monitoring [3, 4], urban planning [5], and resource exploration [6].
@@
CNNs serve as the foundation of many multimodal frameworks due to their ability to capture local spatial patterns through convolutional filters and feature learning [7, 8]. For instance, Hong et al. [9] proposed a CNN-based multimodal learning framework incorporating various fusion strategies, and Wu et al. [10] combined CNNs with a cross-channel reconstruction module. These methods typically use separate CNNs to extract features from each modality and concatenate them along the channel dimension to further learn joint representations. However, CNNs are inherently limited by their receptive fields, which can constrain the modeling of long-range dependencies essential for understanding complex remote sensing scenes.
@@
Transformers mitigate this shortcoming through self-attention mechanisms that capture global context [11]. Xue et al. [12] introduced a deep hierarchical vision transformer leveraging self-attention for spatial features and spectral sequences. Roy et al. [13] proposed a multimodal fusion transformer with a multihead cross-patch attention framework. In addition, numerous transformer-based studies have proposed different modal fusion strategies [14, 15, 16], multiscale feature transformers [17], and global-local structures [18]. Moreover, the Mamba model [19] has gained popularity for capturing long-range dependencies with improved computational efficiency. Building on the Mamba architecture, researchers have developed diffusion-based Mamba [20], spatial-spectral Mamba [21], and multi-scale Mamba [22, 23] variants. In terms of multimodal fusion strategies, Transformer-based methods typically employ global cross-attention mechanisms at the token level, while Mamba-based approaches optimize Mamba blocks to simultaneously perform multimodal fusion and model long-range dependencies.
@@
However, most existing methods are primarily designed to extract features for downstream classification tasks, which limits their ability to capture the complex, high-dimensional, and redundant distributions characteristic of remote sensing data. In addition, challenges remain in modeling complementary diversity features [24, 25] and ensuring robustness to noisy remote sensing data [26]. Diffusion-based two-stage methods have emerged as a promising alternative to other deep learning models for remote sensing image interpretation [27]. By iteratively refining noisy inputs, denoising diffusion probabilistic models (DDPMs) learn robust and noise-reduced representations that capture complex data distributions [28, 29]. For example, Chen et al. [30] employed a diffusion model to extract intrinsic spatial-spectral distribution information from HSI for feature extraction and clustering. Chen et al. [31] proposed SpectralDiff to learn high-dimensional and redundant HSI distributions. In [32] and [33], the multi-step iterative denoising process was used for fusing multi-time-step features for HSI classification. In multimodal scenarios, Zhang et al. [34] integrated diffusion features from HSI into a CNN–Transformer network for classification. Typically, these methods rely on a two-stage pre-training paradigm to train the denoising and feature interpretation networks. To reduce the high computational complexity of heavy pre-training models, it is possible to merge multimodal images into a single pre-training network [35]. However, studies [36, 37, 38] have observed that the modality with more discriminative information tends to dominate the optimization process, leading to the problem of modality imbalance.
@@
Although recent advancements in multimodal remote sensing image classification have yielded significant improvements with respect to more standard approaches, limitations still remain:
@@
DDPMs provide the possibility to capture complex and diverse remote sensing distributions for previous task-oriented models, but few studies have employed diffusion features to guide hierarchical feature extraction and fusion that target both local and global characteristics.
@@
To address these challenges, this paper introduces a balanced diffusion-guided fusion (BDGF) method that employs a pre-training DDPM to guide a multi-branch network in extracting and fusing multimodal features for land-cover classification. Initially, an adaptive modality masking strategy is proposed to pre-train the diffusion model and obtain modality-balanced multimodal diffusion features. Subsequently, these diffusion features guide the CNNs, Mamba, and transformer networks of the multi-branch network through feature fusion, group channel attention, and cross-attention fusion. Finally, a mutual learning approach enhances collaboration among the branches and achieves classification based on the probability entropy and feature similarity of each sub-network. The main contributions of this paper are as follows:
@@
We introduce the BDGF strategy, which leverages multimodal diffusion features to hierarchically guide feature extraction in multi-branch networks based on the global characteristics of features extracted by different branches.
@@
We present a multi-branch network mutual learning strategy that promotes feature complementarity among networks through predicted probability entropy and pairwise feature similarity.
@@ II-A Denoising Diffusion Probabilistic Model
The denoising diffusion probabilistic model (DDPM) [28] is a widely adopted deep generative model [29]. Pre-training a DDPM involves both a forward diffusion process and a reverse diffusion process. Given an image , the noisy image at time can be obtained through a Markov chain:
@@
DDPMs can recover the target distribution through iterative denoising, thereby effectively promoting remote sensing image interpretation tasks [27]. Many studies utilize DDPMs to capture the data distribution of complex images and integrate the extracted features at one [31] or multiple time steps [33] into various tasks during training, such as classification [30, 32], change detection [39, 40], image matching [41], and object detection [42].
In the content of multimodal data analysis, Zhang et al. [34] fused the HSI image features extracted by DDPMs into a downstream multi-branch network to achieve multimodal remote sensing image classification. However, these methods rarely leverage the global guidance provided by diffusion features. Meanwhile, solely extracting diffusion features from HSIs does not satisfy the requirements for multimodal classification. To address these issues, this paper proposes an improved DDPM to obtain modality-balanced multimodal diffusion features. In addition, a series of feature extraction and classification techniques are designed to guide the downstream multi-branch network.
@@ III Methodology
The proposed BDGF is designed to use diffusion distribution to guide the complementarity of diverse features. Fig. 1 illustrates the overall BDGF framework. In the pre-training phase, the denoising network generates modality-balanced diffusion features using an adaptive modality masking strategy. During training, these features hierarchically guide the extraction and fusion of information within the multi-branch network. Finally, the mutual learning module enhances the alignment of multimodal features.
The following subsections provide a detailed explanation of each module within the BDGF framework, highlighting the related functionalities and interactions.
@@ III-A Adaptive Modality Masking-Based DDPMs
Inspired by [36, 38], we aim to prevent DDPMs from over-focusing on spectral images. Unlike reconstruction tasks that add large blocks of masks from a spatial or spectral perspective [46, 47], our strategy employs a sample mask and an image structure mask for spectral images . The sample mask reduces the proportion of the dominant modal in the batch, while randomly masks a certain proportion of the image using a minimal block. Assuming that represents the model iteration process, the mask generation can be expressed as:
where is Bernoulli distribution. This distribution easily allows unbiased, element-wise random masking based on a single parameter, the iteration process. Based on (5)-(7), we can merge the SAR image and spectral image along the channel dimension. The noisy image at time t can be expressed as:
Given and as SAR Gamma shape and distribution respectively, the transfer probability of forward diffusion process can be expressed as:
Correspondingly, the transfer probability of reverse diffusion process can be expressed as:
where and are the Gamma shape and scale, respectively. and are mean and variance similar to (3). The noise predicted by the model is .
A limitation of merging along the channel dimension is that it reduces information exchange between modalities. To mitigate this issue, random channel exchange is incorporated into the down-sampling process of the denoising network, as shown in Fig. 1. This method randomly swaps channels in spectral and SAR features, thereby promoting the fusion of multimodal features across multiple time steps and iterations. Fig. 3 visualizes the distribution effect of the features obtained by the proposed method on the LCZ HK dataset. The execution details are the same as in Section III-C. Across categories the adaptive multimodality masking (a) produces more uniform, consolidated class clusters and reduces dominance and fragmentation seen in the original model. For example, class C6 forms a tight, isolated cluster in (a) but is scattered and intermixed with other classes in (b).
@@ III-B Diffusion Features Guidance
To leverage the multimodal data distribution extracted through the diffusion process, the diffusion features guide the extraction of diverse features in accordance with the characteristics of each branch. CNN-based architectures capture local features using small convolution kernels, whereas transformers and Mamba models extract global features via attention mechanisms and state space models. Mamba is particularly well-suited for tasks involving long sequence features [48]. Section III-C presents a detailed visualization analysis of the complementarity of multi-branch network features.
@@ III-B1 CNN-Based Network
where and denote three-dimensional and two-dimensional convolution operations, respectively. The intermediate feature vectors , , and are produced during the network’s processing, and , , and are trainable scalar parameters. The output of the feature-fusion-based CNN module is given by:
@@ III-C Mutual Learning Module
The mutual learning module promotes collaboration among sub-networks and enhances the fusion of diverse features [14]. The experiments in Section III-C visualize the effect of this strategy. As illustrated in Fig. 7, this module uses KL divergence to align the entropy and feature similarity of paired networks.
@@ IV-A4 LCZ HK dataset (MSI+SAR)
In our experiments, we employ four metrics to quantitatively assess classification performance: class-specific accuracy, overall accuracy (OA), average accuracy (AA), and the kappa coefficient (Kappa). The experiments are implemented using PyTorch and executed on an NVIDIA GeForce RTX 3090 (24 GB). Optimization is performed using the Adam algorithm with a learning rate of , modulated by a MultiStepLR scheduler with a decay factor of 0.5. The patch size is 9 and the dimension of embedding features is 64. For comparative experiments, all HSI and SAR datasets use 100 labeled samples per class, while the MSI and SAR dataset uses 50 labeled samples per class.
@@ IV-C Feature Complementarity Visualization
To verify that our multi-branch architecture learns complementary representations, we visualize the per-branch features immediately before the final fusion layer. Fig. 9. shows features extracted from the LCZ HK dataset (13 classes, C0–C12, 50 samples per class) projected onto two dimensions via t-distributed stochastic neighbor embedding (t-SNE). The t-SNE projection is generated from the first 50 principal components of the data, using a perplexity of 30 and 1000 iterations to ensure reproducible results.
@@ IV-D Classification Results
Furthermore, we selected four methods that focus on multi-branch network structures. Fusion-HCT and MACN integrate CNNs and transformers to capture both local and global features, introducing innovative attention mechanisms for multimodal feature fusion. NCGLF enhances CNN and transformer structures with structural information learning and invertible neural networks. UACL proposes a contrastive learning strategy to access reliable multimodal samples. In addition, two recent multimodal learning methods based on the multi-scale Mamba structure are included for comparison. HLMamba constructs a multimodal Mamba fusion module and introduces a gradient joint algorithm to enhance modal information, while MSFMamba employs spatial, spectral, and fused Mamba branches with a large effective receptive field to achieve multi-scale feature fusion.
The performance of these classification methods is summarized in Tables III–VI. For a visual comparison, the respective segmentation maps of several methods are presented in Figs. 10-13 To further verify the effectiveness of the model on different training samples, we conducted experiments with 60, 80, 120, and 140 labeled samples per class for HSI and SAR classification, and 30, 40, 60, and 70 labeled samples per class for MSI and SAR classification. The performance of each method under these settings is illustrated in Fig. 14 Based on these outcomes, the following conclusions can be drawn:

### v2

@@ Abstract
Deep learning-based techniques for the analysis of multimodal remote sensing data have become popular due to their ability to effectively integrate spatial, spectral, and structural information from different sensors. Recently, denoising diffusion probabilistic models (DDPMs) have attracted attention in the remote sensing community due to their powerful ability to capture robust and complex spatial-spectral distributions. However, pre-training multimodal DDPMs may result in modality imbalance, and effectively leveraging diffusion features to guide complementary diversity feature extraction remains an open question. To address these issues, this paper proposes a balanced diffusion-guided fusion framework that leverages multimodal diffusion features to guide a multi-branch network for land-cover classification. Specifically, we propose an adaptive modality masking strategy to encourage the DDPMs to obtain a modality-balanced rather than spectral image-dominated data distribution. Subsequently, these diffusion features hierarchically guide feature extraction among CNN, Mamba, and transformer networks by integrating feature fusion, group channel attention, and cross-attention mechanisms. Finally, a mutual learning strategy is developed to enhance inter-branch collaboration by aligning the probability entropy and feature similarity of individual subnetworks. Extensive experiments on four multimodal remote sensing datasets demonstrate that the proposed method achieves superior classification performance. The code is available at https://github.com/HaoLiu-XDU/BDGF.
@@
Multimodal remote sensing data typically comprises images from various sensors, including hyperspectral images (HSI), synthetic aperture radar (SAR), and light detection and ranging. By fusing complementary spatial, spectral, and structural information, multimodal data facilitates an accurate classification [1], thereby supporting applications in environmental monitoring [2, 3], urban planning [4], and resource exploration [5].
@@
CNNs serve as the foundation of many multimodal frameworks due to their ability to capture local spatial patterns through convolutional filters and feature learning [6, 7]. For instance, Hong et al. [8] proposed a CNN-based framework with various fusion strategies, and Wu et al. [9] combined CNNs with a cross-channel reconstruction module. These methods typically use separate CNNs to extract features from each modality and concatenate them along the channel dimension to further learn joint representations. However, CNNs are inherently limited by their receptive fields, which can constrain the modeling of long-range dependencies essential for understanding complex remote sensing scenes.
@@
Transformers mitigate this shortcoming through self-attention mechanisms that capture global context [10]. Xue et al. [11] introduced a deep hierarchical vision transformer leveraging self-attention for spatial features and spectral sequences. Roy et al. [12] proposed a multimodal fusion transformer with a multihead cross-patch attention framework. In addition, numerous transformer-based studies have proposed different fusion strategies [13, 14], multiscale feature transformers [15], and global-local structures [16]. Moreover, the Mamba model [17] has gained popularity for capturing long-range dependencies with improved computational efficiency. Building on the Mamba architecture, researchers have developed diffusion-based Mamba [18], spatial-spectral Mamba [19], and multi-scale Mamba [20, 21] variants. In terms of multimodal fusion strategies, Transformer-based methods typically employ global cross-attention mechanisms at the token level, while Mamba-based approaches optimize Mamba blocks to simultaneously perform multimodal fusion and model long-range dependencies. Adaptively merging these local, long-range, and sequence-aware representations can leverage their strengths and compensate for their weaknesses [22, 23, 24].
However, most existing methods are primarily designed to extract features for downstream classification task, which limits their ability to capture the complex, high-dimensional, and redundant distributions characteristic of remote sensing data. In addition, challenges remain in modeling complementary diversity features [25, 26] and ensuring robustness to noisy remote sensing data [27].
@@
Diffusion-based two-stage methods have emerged as a promising alternative to other deep learning models for remote sensing image interpretation [28]. By iteratively refining noisy inputs, denoising diffusion probabilistic models (DDPMs) learn robust and noise-reduced representations that capture complex data distributions [29, 30]. For example, Chen et al. [31] and Chen et al. [32] both employed a diffusion model to extract high-dimensional and redundant distribution from HSI. In [33] and [34], the multi-step iterative denoising process was used for fusing multi-time-step features for HSI classification. In multimodal scenarios, Zhang et al. [35] only integrated HSI diffusion features instead of multimodal information into the downstream network, which is essentially hyperspectral feature extraction and classification. Typically, these methods rely on a two-stage pre-training paradigm to train the denoising and feature interpretation networks. To reduce the high computational complexity of heavy pre-training models, it is possible to merge multimodal images into a single pre-training network [36]. However, studies [37, 38, 39] have observed that the modality with more discriminative information tends to dominate the optimization process, leading to the problem of modality imbalance.
@@
Although recent advancements have yielded significant improvements, limitations still remain:
@@
DDPMs capture complex remote sensing distributions for previous task-oriented models. However, few studies have employed global and noise-robust diffusion information to guide the extraction of various characteristics, including local, long-range, and sequence-aware features.
@@
Fig. 1 compares the workflows and illustrates the motivation behind our approach. To address these challenges, this paper introduces a balanced diffusion-guided fusion (BDGF) method that employs a pre-training DDPM to guide a multi-branch network for multimodal land-cover classification. Initially, an adaptive modality masking strategy is proposed to pre-train the diffusion model and obtain modality-balanced multimodal diffusion features. Subsequently, these diffusion features guide the CNNs, Mamba, and transformer networks of the multi-branch network through feature fusion, group channel attention, and cross-attention fusion. Finally, a mutual learning approach dynamically enhances collaboration among the branches and achieves classification based on the probability entropy and feature similarity of each sub-network. The main contributions of this paper are as follows:
@@
We introduce a diffusion-guided fusion strategy that leverages multimodal diffusion information to hierarchically guide the fusion of diverse features based on the global characteristics of different branches.
@@
We present a multi-branch network mutual learning strategy that dynamically promotes feature complementarity among networks through predicted probability entropy and pairwise similarity.
@@ II-A Denoising Diffusion Probabilistic Model
The DDPM [29] is a widely adopted deep generative model [30]. Pre-training a DDPM involves both a forward diffusion process and a reverse diffusion process. Given an image , the noisy image at time can be obtained through a Markov chain:
@@
DDPMs can recover the target distribution through iterative denoising, thereby effectively promoting remote sensing image interpretation tasks [28]. Many studies utilize DDPMs to capture the data distribution of complex images and integrate the extracted features at one [32] or multiple time steps [34] into various tasks during training, such as classification [31, 33], change detection [40, 41], image matching [42], and object detection [43]. In the content of multimodal data analysis, Zhang et al. [35] fused the HSI image features extracted by DDPMs into a downstream multi-branch network. Du et al. [18] integrated diffusion feature into mamba structure for semantic segmentation.
However, these methods rarely leverage the global guidance provided by diffusion features. Meanwhile, solely extracting diffusion features from HSIs does not satisfy the requirements for multimodal classification. To address these issues, this paper proposes an improved DDPM to obtain modality-balanced multimodal diffusion features. In addition, a series of feature extraction and classification techniques are designed to guide the downstream multi-branch network.
@@ III Methodology
The proposed BDGF is designed to use diffusion distribution to guide the complementarity of diverse features. Fig. 2 illustrates the overall BDGF framework. In the pre-training phase, the denoising network generates modality-balanced diffusion features using an adaptive modality masking strategy. During training, these features hierarchically guide the extraction and fusion of information within the multi-branch network. Finally, the mutual learning module enhances the alignment of multimodal features. The following subsections provide a detailed explanation of each module within the BDGF framework.
@@ III-A Adaptive Modality Masking-Based DDPMs
Inspired by [37, 39], we aim to prevent DDPMs from over-focusing on spectral images. Unlike reconstruction tasks that add large blocks of masks from a spatial or spectral perspective [46, 47], our strategy dynamically employs a sample mask and a structure mask for spectral images . The sample mask reduces the proportion of the dominant modality in the batch, while randomly masks a certain proportion of the image using a minimal block. To dynamically suppress the dominant modality, the mask ratio is continuously increased during the iterative process. Assuming that represents the training progress, the mask generation can be expressed as:
where is Bernoulli distribution and is an exponential operation. This soft distribution allows unbiased and element-wise random masking based on the iteration progress. Based on (5)-(7), we add multiplicative speckle noise to the SAR image and Gaussian noise to the spectral image . Then, we merge them along the channel dimension. (1) can be transformed into:
Given and as SAR Gamma shape and distribution respectively, the transfer probability of forward diffusion process (2) can be expressed as:
Correspondingly, the transfer probability of reverse diffusion process (3) can be expressed as:
where and are the Gamma shape and scale, respectively. and are mean and variance similar to (3). The denoising network updates its parameters via gradient descent. The pseudo-code is shown in Algorithm 1.
Fig. 4 visualizes the distribution effect of the features obtained by the proposed method on the LCZ HK dataset. The t-SNE execution details are the same as in Section IV-C. Across categories the adaptive multimodality masking (a) produces more uniform, consolidated class clusters and reduces dominance and fragmentation seen in the original model. For example, class C6 forms a tight, isolated cluster in (a) but is scattered and intermixed with other classes in (b).
@@ III-B Diffusion Features Guidance
To leverage the multimodal data distribution extracted through the diffusion process, the diffusion features guide the extraction of diverse features in accordance with the characteristics of each branch. CNN-based architectures capture local features using small convolution kernels, whereas transformers and Mamba models extract global features via attention mechanisms and state space models. Mamba is particularly well-suited for tasks involving long sequence features [48]. Section IV-C presents a detailed visualization analysis of the complementarity of multi-branch network features.
@@ III-B1 CNN-Based Network
where and denote three-dimensional and two-dimensional convolution operations, respectively. The intermediate feature vectors , , and are produced during the network’s processing, and , , and are trainable scalar parameters. The output of the feature fusion-based CNN module is given by:
@@ III-C Mutual Learning Module
The mutual learning module promotes collaboration among sub-networks and enhances the fusion of diverse features [13]. The experiments in Section IV-C visualize the effect of this strategy. As illustrated in Fig. 8, this module uses KL divergence to align the entropy and feature similarity of paired networks.
@@ IV-A4 LCZ HK dataset (MSI+SAR)
In our experiments, we employ four metrics to quantitatively assess classification performance: class-specific accuracy, overall accuracy (OA), average accuracy (AA), and the kappa coefficient (Kappa). The experiments are implemented using PyTorch and executed on an NVIDIA GeForce RTX 3090 (24 GB). The denoising network follows the U-Net structure of SpectralDiff [32] but adds our masking strategy. Optimization is performed using the Adam algorithm with a learning rate of , modulated by a MultiStepLR scheduler with a decay factor of 0.5. The patch size is 9 and the dimension of embedding features is 64. The implementation parameter details are in the publicly available code to ensure reproducibility 111https://github.com/HaoLiu-XDU/BDGF. All the comparison methods are executed under the same configurations. HSI and SAR datasets use 100 labeled samples per class, while the MSI and SAR dataset uses 50 labeled samples per class.
@@ IV-C Feature Complementarity Visualization
To verify that our multi-branch architecture learns complementary representations, we visualize the per-branch features immediately before the final fusion layer. Fig. 10 shows features extracted from the LCZ HK dataset (13 classes, C0–C12, 50 samples per class) projected onto two dimensions via t-distributed stochastic neighbor embedding (t-SNE). The t-SNE projection is generated from the first 50 principal components of the data, using a perplexity of 30 and 1000 iterations to ensure reproducible results.
@@ IV-D Classification Results
Furthermore, we selected four methods that focus on multi-branch network structures. Fusion-HCT and MACN integrate CNNs and transformers to capture both local and global features, introducing innovative attention mechanisms for multimodal feature fusion. NCGLF enhances CNN and transformer structures with structural information learning and invertible neural networks. UACL proposes a contrastive learning strategy to access reliable multimodal samples. In addition, two recent multimodal learning methods based on the multi-scale Mamba structure are included for comparison. HLMamba constructs a multimodal Mamba fusion module and introduces a gradient joint algorithm to enhance modality information, while MSFMamba employs spatial, spectral, and fused Mamba branches with a large effective receptive field to achieve multi-scale feature fusion.
The performance of these classification methods is summarized in Tables III–VI. For a visual comparison, the respective segmentation maps of several methods are presented in Figs. 11-14. To further verify the effectiveness of the model on different training samples, we conducted experiments with 60, 80, 120, and 140 labeled samples per class for HSI and SAR classification, and 30, 40, 60, and 70 labeled samples per class for MSI and SAR classification. The performance of each method under these settings is illustrated in Fig. 15. Based on these outcomes, the following conclusions can be drawn:
@@ IV-E Analysis of the Computational Complexity
In this section, we evaluate the computational complexity of all models in terms of GFLOPs and number of parameters (millions). Table VII presents these metrics for the four datasets considered. Note that SS-MAE and BDGF employ a pre-training paradigm, and (P) and (T) denote pre-training and training phases, respectively. Under unsupervised learning, larger models generally generalize better. Pre-training methods incur higher computational costs than other approaches, and adopting a dual-branch structure for the pre-training autoencoder and diffusion model further increases complexity. SS-MAE and BDGF mitigate this issue by merging multimodal images along the channel dimension. Their superior classification results across the four datasets confirm the effectiveness of pre-training. For the proposed BDGF, the parameter count arises from integrating multi-branch sub-networks, while the FLOPs reflect the diffusion model’s pre-training. Although BDGF exhibits higher computational complexity, it remains within acceptable limits and achieves the best classification performance.
IV-F Analysis of the Transferability